# Foundation Models for Causal Inference via Prior-Data Fitted Networks

### v2

@@ 1 Introduction
Causal inference is a cornerstone of empirical research in disciplines such as economics (Angrist, 1990; Imbens & Angrist, 1994), medicine (Feuerriegel et al., 2024; Weberpals et al., 2025), and marketing (Varian, 2016). It enables the estimation of causal effects from observational and randomized data, which is essential for reliable decision-making (Kern et al., 2025). In personalized medicine, for instance, it supports identifying the most effective treatment by predicting patient outcomes under different therapeutic options.
In recent years, machine learning, and especially deep learning methods, have gained significant traction in causal inference (Curth & van der Schaar, 2021; Ma et al., 2025; 2024; Melnychuk et al., 2022; Schweisthal et al., 2023; Shalit et al., 2017a; Shi et al., 2019)). These methods offer several advantages for causal effect estimation in practice, including the ability to handle large-scale and high-dimensional datasets with complex confounding structures and to model heterogeneity of causal effects (Feuerriegel et al., 2025). However, most existing approaches require retraining a model for each new dataset. To this end, existing approaches lack the flexibility to perform inference for new datasets without additional retraining, which limits their practicality in real-world settings.
@@
Our contributions [^1] Code is available at https://github.com/yccm/CausalFM. are: (1) We formalize the constructions of priors based on structural causal models (SCMs) for Bayesian causal inference and derive necessary conditions for their validity. (2) We propose an explicit CausalFM prior based on Bayesian neural networks that are compatible with the structure of the causal inference problem at hand. We also propose a learning algorithm to train PFNs for causal inference problems that leverages our CausalFM prior to simulate counterfactuals to mitigate the fundamental problem of causal inference. (3) We propose concrete instantiations of our framework by training PFNs for estimating conditional average treatment effects (CATEs) in different causal inference settings. We show empirically that CausalFM performs competitively and outperforms current state-of-the-art CATE estimators on a variety of benchmarks.
@@ 2 Related Work
Amortized causal inference. Several recent papers pre-train large neural networks on synthetic data so that they can solve causal tasks via in-context learning. Examples include causal discovery (Mahajan et al., 2025), ATE estimation under unconfoundedness (Zhang et al., 2024), zero-shot- and few-shot learning (Nilforoshan et al., 2023; Iwata & Chikahara, 2023), and reinforcement-learning (Lee et al., 2023). These methods validate the feasibility of treating causal inference as an in-context learning problem but remain restricted to specific causal inference settings, which typically do not allow accommodating unobserved confounding.
@@ 3.2 Task: causal inference
As an illustrative example, we consider a standard causal inference setting, called backdoor adjustment, where the data comprise $(X,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$, where $X$ are patient covariates, $A$ is a treatment, and $Y$ is an outcome of interest (van der Laan & Rubin, 2006). For example, in medicine, $X$ may contain treatment history or demographic attributes, $A$ may be a medical treatment, and $Y$ a health outcome. Following the potential outcome framework (Rubin, 1974), let $Y(a)$ denote the outcome that would be realized under the treatment $A=a$. The interventional distribution is thus over $(X,A,Y(1)-Y(0))\sim\mathbb{P}_{\mathrm{int}}$, and a common target functional is the conditional average treatment effect (CATE) $Q(x)\;=\;\mathbb{E}\bigl[Y(1)-Y(0)\mid X=x\bigr]$ (Wager & Athey, 2018). The CATE quantifies the expected benefit of providing treatment given the patient’s covariates.
Identifiability. To estimate CATE from observational data, we need to impose identifiability assumptions, which link the observational and the interventional distributions and allow us to express $Q$ as a functional of the observational distribution (Rosenbaum & Rubin, 1983). These are (i) consistency: $Y(A)=Y$, (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)>0$, and (iii) Unconfoundedness: $Y(1),Y(0)\perp\!\!\!\perp A\mid X$ in $\mathbb{P}_{\mathrm{int}}$.
@@ 3.2.1 Running examples
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy the following conditions (Newey & Powell, 2003): (i): Additive structural equation: $Y=f(X,A)+g(X,U)$, with (unknown) functions $f$ and $g$ and unobserved confounder $U$, implying that $Y$ does not directly depend on $Z$; (ii) Independence: $U\perp\!\!\!\perp Z\mid X$; (iii) Relevance: $\mathbb{P}_{\mathrm{obs}}(A\mid X=x,Z=z)>0$ is non-constant in $z$; and (iv) Completeness: For every measurable $g$, if $\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[f(x,A)\mid X=x,Z=z]=0$ for all $z$, then $f(x,A)=0$ almost surely in $\mathbb{P}_{\mathrm{obs}}$. Then, the CATE can be shown to be identified via an integral equation (Newey & Powell, 2003; Hartford et al., 2017).
@@ 4.1 Introducing SCM-based priors
However, the above approach has drawbacks: (i) It requires knowledge of the identification formula $\bar{Q}$, which must be determined on a case-by-case basis depending on the causal inference setting $\mathcal{C}$ at hand. This can be tedious or even hard to compute in practice. For example, the IV setting from Example 3 requires solving integral equations to compute $\bar{Q}$ (Newey & Powell, 2003). (ii) Constructing a prior for $\mathbb{P}_{\mathrm{obs}}$ makes it harder control the distribution of the causal query $Q$ directly. It has been shown in the literature that this can lead to prior misspecification for Bayesian causal inference or slowly converging posterior distributions (Linero & Antonelli, 2022).
@@ 4.2 Well-specified priors
Theorem 4.3.
Let $\mathcal{Z}$ be the set of all identifiability-violating SCMs $\mathcal{S}_{0}$ that satisfy $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}}\in\mathcal{P}_{\mathrm{obs}}$ and $Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{0}}\bigr)\neq\bar{Q}\!\bigl(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}}\bigr)$. Assume that $Q$ is a linear functional (e.g., the CATE) and that $\int_{\mathcal{Z}}Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\,\Pi(\mathcal{S})\,\mathop{}\!\mathrm{d}\mathcal{S}\neq\bar{Q}\!\bigl(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{0}}\bigr)$ (non-identifiability doesn’t cancel out). Then, if $\Pi(\mathcal{S})$ is well-specified for $\mathcal{C}$, it follows that $\Pi(\mathcal{S}\in\mathcal{Z})=0$.
@@ 5.1 Training algorithm
6 Experiments
Evaluation metrics. We report the precision in estimating heterogeneous effects (PEHE) (Curth & van der Schaar, 2021; Hill, 2011), defined as the root mean squared deviation between predicted and ground-truth CATE, to evaluate the model performance on the CATE estimation task.
6.1 Evaluation for standard CATE setting
Baselines for standard CATE estimation. We consider a broad range of state-of-the-art methods for the conditional treatment effect estimation from the literature: (1) S-learner (Künzel et al., 2019): the S-learner is a model-agnostic learner that trains a single regression model by concatenating the covariate and the treatment as input; (2) T-learner (Künzel et al., 2019): the T-learner is a model-agnostic learner that trains separate regression models for treated and control groups; (3) X-learner (Künzel et al., 2019): builds upon the T-learner by first imputing individual treatment effects in each group and then fitting models to these pseudo-effects; (4) TARNet (Shalit et al., 2017b): using representation learning to extract features of covariates and train separate branches for treated and control groups with regularization; (5) DR-learner (Kennedy, 2023b): generates pseudo-outcomes based on the doubly-robust AIPW estimator; (6) RA-learner (Curth & van der Schaar, 2021): uses a regression-adjusted pseudo-outcome in the second stage. We also include two PFN-based foundation models for treatment effect estimation: (7) CausalPFN (Balazadeh et al., 2025) and (8) DoPFN (Robertson et al., 2025). Further implementation details are in Appendix D.3.
@@ 6.1 Evaluation for standard CATE setting
RA-learner (Curth & van der Schaar, 2021) 0.609 $\pm$ 0.13 0.652 $\pm$ 0.15
@@
Results on standard CATE estimation. We benchmark our model on ten synthetic datasets generated under diverse mechanisms, with implementation details in Appendix D. In addition, we evaluate on a semi-synthetic version of the Jobs dataset (Smith & Todd, 2005), derived from the widely used LaLonde study (LaLonde, 1986). Here, we generate outcomes to create a semi-synthetic dataset and allow for evaluation against ground-truth.
@@
MRIV (Frauen & Feuerriegel, 2022) 0.688 $\pm$ 0.21 0.641 $\pm$ 0.24
@@ 6.2 Evaluation for IV setting
Baselines for IV setting. We benchmark against a broad set of state-of-the-art IV methods for treatment effect estimation: (1) KIV (Singh et al., 2019): a nonlinear extension of two-stage least squares using kernel ridge regression with feature maps; (2) DFIV (Xu et al., 2021): extends KIV by parameterizing feature maps with neural networks trained iteratively; (3) DeepIV (Hartford et al., 2017): a two-stage neural approach, first estimating the treatment distribution and then solving a counterfactual prediction task; (4) DeepGMM (Bennett et al., 2019): formulates IV estimation as a minimax game based on the generalized method of moments, solved via adversarial training; (5) DMLIV (Syrgkanis et al., 2019): a double machine learning framework that estimates nuisance functions and learns the CATE by orthogonalized regression; (6) DRIV (Syrgkanis et al., 2019): a meta-learner combining DMLIV with doubly robust pseudo-outcomes for improved stability; and (7) MRIV (Frauen & Feuerriegel, 2022): a multiply robust framework for binary IVs that directly estimates CATE via pseudo–outcome regression. For foundation model baselines, as CausalPFN (Balazadeh et al., 2025) is only for back-door adjustment, we include DoPFN (Robertson et al., 2025).
@@ 6.3 Front-door adjustment
Baselines for front-door adjustment setting. In contrast to the standard CATE or IV settings, there are few established baselines for the front-door case. Identification in this setting is enabled through Pearl’s front-door formula (Pearl, 2009). The natural baseline is therefore the plug-in front-door learner, which estimates the necessary nuisance components, i.e., $P(M\mid A,X)$, $P(A\mid X)$, and $\mathbb{E}[Y\mid M,X]$ and substitutes them into the identification formula to recover causal quantities. To assess the role of model flexibility in estimating these nuisance functions, we implement the plug-in learner with different regression methods, including linear regression, Random Forests, and neural networks.
Results for front-door adjustment setting. We additionally evaluate our model under the front-door adjustment setting in Table 6.2. Table 6.2 reports the averaged PEHE across datasets. We observe that CausalFM achieves competitive CATE estimation. Importantly, these results hold without requiring model retraining for our model, demonstrating the flexibility of our method to perform causal inference in the front-door adjustment setting.
7 Conclusion and Future Work
This paper introduces CausalFM, a framework for training foundation models for causal inference using prior-data fitted networks (PFNs). Our framework is capable of handling a broad range of causal inference settings, including standard CATE estimation, instrumental variables, and front-door adjustment. Looking ahead, we plan to further scale CausalFM to larger and more diverse datasets to assess its generalizability and practical utility. In doing so, our goal is to establish robust foundation models that can serve as a unifying approach for a wide spectrum of causal inference tasks and datasets.
Acknowledgments
This paper is supported by the DAAD program “Konrad Zuse Schools of Excellence in Artificial Intelligence”, sponsored by the Federal Ministry of Education and Research. Additionally, this work has been supported by the German Federal Ministry of Education and Research (Grant: 01IS24082).

### v3

@@ 1 Introduction
Causal inference is a cornerstone of empirical research in disciplines such as economics (Angrist, 1990; Imbens and Angrist, 1994), medicine (Feuerriegel et al., 2024; Weberpals et al., 2025), and marketing (Varian, 2016). It enables the estimation of causal effects from observational and randomized data, which is essential for reliable decision-making (Kern et al., 2025). In personalized medicine, for instance, it supports identifying the most effective treatment by predicting patient outcomes under different therapeutic options.
In recent years, machine learning, and especially deep learning methods, have gained significant traction in causal inference (Curth and van der Schaar, 2021a; Ma et al., 2025; 2024; Melnychuk et al., 2022; Schweisthal et al., 2023; Shalit et al., 2017a; Shi et al., 2019)). These methods offer several advantages for causal effect estimation in practice, including the ability to handle large-scale and high-dimensional datasets with complex confounding structures and to model heterogeneity of causal effects (Feuerriegel et al., 2025). However, most existing approaches require retraining a model for each new dataset. To this end, existing approaches lack the flexibility to perform inference for new datasets without additional retraining, which limits their practicality in real-world settings.
@@
Our contributions [^1] Our CausalFM Toolkit is available at https://github.com/yccm/CausalFM-toolkit; Toolkit docs are at https://causalfm-toolkit.readthedocs.io; Project code is available at https://github.com/yccm/CausalFM. are: (1) We formalize the constructions of priors based on structural causal models (SCMs) for Bayesian causal inference and derive necessary conditions for their validity. (2) We propose an explicit CausalFM prior based on Bayesian neural networks that are compatible with the structure of the causal inference problem at hand. We also propose a learning algorithm to train PFNs for causal inference problems that leverages our CausalFM prior to simulate counterfactuals to mitigate the fundamental problem of causal inference. (3) We propose concrete instantiations of our framework by training PFNs for estimating conditional average treatment effects (CATEs) in different causal inference settings. We show empirically that CausalFM performs competitively and outperforms current state-of-the-art CATE estimators on a variety of benchmarks.
@@ 2 Related Work
Amortized causal inference. Several recent papers pre-train large neural networks on synthetic data so that they can solve causal tasks via in-context learning. Examples include causal discovery (Mahajan et al., 2025), ATE estimation under unconfoundedness (Zhang et al., 2024), zero-shot- and few-shot learning (Nilforoshan et al., 2023; Iwata and Chikahara, 2023), and reinforcement-learning (Lee et al., 2023). These methods validate the feasibility of treating causal inference as an in-context learning problem but remain restricted to specific causal inference settings, which typically do not allow accommodating unobserved confounding.
@@
Finally, Dhir et al. (2025) propose a framework for amortized estimation of interventional distributions using neural processes (Garnelo et al., 2018). Our work differs in that we focus on inference on identified causal quantities without knowledge of the full causal graph (e.g., using clusters of confounders).
@@ 3.2 Task: causal inference
As an illustrative example, we consider a standard causal inference setting, called backdoor adjustment, where the data comprise $(X,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$, where $X$ are patient covariates, $A$ is a treatment, and $Y$ is an outcome of interest (van der Laan and Rubin, 2006). For example, in medicine, $X$ may contain treatment history or demographic attributes, $A$ may be a medical treatment, and $Y$ a health outcome. Following the potential outcome framework (Rubin, 1974), let $Y(a)$ denote the outcome that would be realized under the treatment $A=a$. The interventional distribution is thus over $(X,A,Y(1)-Y(0))\sim\mathbb{P}_{\mathrm{int}}$, and a common target functional is the conditional average treatment effect (CATE) $Q(x)\;=\;\mathbb{E}\bigl[Y(1)-Y(0)\mid X=x\bigr]$ (Wager and Athey, 2018). The CATE quantifies the expected benefit of providing treatment given the patient’s covariates.
Identifiability. To estimate CATE from observational data, we need to impose identifiability assumptions, which link the observational and the interventional distributions and allow us to express $Q$ as a functional of the observational distribution (Rosenbaum and Rubin, 1983). These are (i) consistency: $Y(A)=Y$, (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)>0$, and (iii) Unconfoundedness: $Y(1),Y(0)\perp\!\!\!\perp A\mid X$ in $\mathbb{P}_{\mathrm{int}}$.
@@ 3.2.1 Running examples
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy the following conditions (Newey and Powell, 2003): (i): Additive structural equation: $Y=f(X,A)+g(X,U)$, with (unknown) functions $f$ and $g$ and unobserved confounder $U$, implying that $Y$ does not directly depend on $Z$; (ii) Independence: $U\perp\!\!\!\perp Z\mid X$; (iii) Relevance: $\mathbb{P}_{\mathrm{obs}}(A\mid X=x,Z=z)>0$ is non-constant in $z$; and (iv) Completeness: For every measurable $g$, if $\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[f(x,A)\mid X=x,Z=z]=0$ for all $z$, then $f(x,A)=0$ almost surely in $\mathbb{P}_{\mathrm{obs}}$. Then, the CATE can be shown to be identified via an integral equation (Newey and Powell, 2003; Hartford et al., 2017).
@@ 4.1 Introducing SCM-based priors
However, the above approach has drawbacks: (i) It requires knowledge of the identification formula $\bar{Q}$, which must be determined on a case-by-case basis depending on the causal inference setting $\mathcal{C}$ at hand. This can be tedious or even hard to compute in practice. For example, the IV setting from Example 3 requires solving integral equations to compute $\bar{Q}$ (Newey and Powell, 2003). (ii) Constructing a prior for $\mathbb{P}_{\mathrm{obs}}$ makes it harder control the distribution of the causal query $Q$ directly. It has been shown in the literature that this can lead to prior misspecification for Bayesian causal inference or slowly converging posterior distributions (Linero and Antonelli, 2022).
@@ 4.2 Well-specified priors
Theorem 4.3 (Identifiability rules out prior mass on violating SCMs).
Assume the causal inference setting $\mathcal{C}$ is identifiable for the causal query $Q$ with corresponding identified functional $\bar{Q}$. Let
$\mathcal{Z}:=\Bigl\{\mathcal{S}:\ \mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}\in\mathcal{P}_{\mathrm{obs}}\ \text{ and }\ Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\neq\bar{Q}\!\bigl(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}\bigr)\Bigr\}$
denote the set of identifiability-violating SCMs, and for any $P_{\mathrm{obs}}\in\mathcal{P}_{\mathrm{obs}}$ define the observational equivalence class
$\mathcal{E}(P_{\mathrm{obs}}):=\{\mathcal{S}:\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}=P_{\mathrm{obs}}\}.$
Assume $Q$ is a linear functional of $\mathbb{P}_{\mathrm{int}}$ and that non-identifiability does not cancel out in prior expectation, i.e., for every $P_{\mathrm{obs}}\in\mathcal{P}_{\mathrm{obs}}$ with $\Pi\!\bigl(\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})\bigr)>0$,
$\int_{\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})}Q\!\bigl(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\bigr)\,\Pi(\mathop{}\!\mathrm{d}\mathcal{S})\;\neq\;\bar{Q}(P_{\mathrm{obs}})\,\Pi\!\bigl(\mathcal{Z}\cap\mathcal{E}(P_{\mathrm{obs}})\bigr).$
Then any prior $\Pi(\mathcal{S})$ that is well-specified for $\mathcal{C}$ must satisfy $\Pi(\mathcal{Z})=0$.
@@ 4.3 Constructing SCM-based priors
Notes on identifiability. Our framework follows established causal inference philosophy and separates identifiability from estimation (Pearl, 2009): the identifiability step (==choosing the causal setting) requires careful modeling and usage of domain knowledge, while the estimation step can be handed over to our CausalFM. If practitioners suspect identifiability assumptions may be violated, we recommend performing causal sensitivity analysis (Dorn and Guo, 2022; Frauen et al., 2023) to assess the extent of potential violations.
@@ 5.1 Training algorithm
RA-learner (Curth and van der Schaar, 2021b) 0.609 $\pm$ 0.13 0.652 $\pm$ 0.15
@@
6 Experiments
Evaluation metrics. We report the precision in estimating heterogeneous effects (PEHE) (Curth and van der Schaar, 2021b; Hill, 2011), defined as the root mean squared deviation between predicted and ground-truth CATE, to evaluate the model performance on the CATE estimation task.
6.1 Evaluation for standard CATE setting
Baselines for standard CATE estimation. We consider a broad range of state-of-the-art methods for the conditional treatment effect estimation from the literature: (1) S-learner (Künzel et al., 2019): the S-learner is a model-agnostic learner that trains a single regression model by concatenating the covariate and the treatment as input; (2) T-learner (Künzel et al., 2019): the T-learner is a model-agnostic learner that trains separate regression models for treated and control groups; (3) X-learner (Künzel et al., 2019): builds upon the T-learner by first imputing individual treatment effects in each group and then fitting models to these pseudo-effects; (4) TARNet (Shalit et al., 2017b): using representation learning to extract features of covariates and train separate branches for treated and control groups with regularization; (5) DR-learner (Kennedy, 2023b): generates pseudo-outcomes based on the doubly-robust AIPW estimator; (6) RA-learner (Curth and van der Schaar, 2021b): uses a regression-adjusted pseudo-outcome in the second stage. We also include two PFN-based foundation models for treatment effect estimation: (7) CausalPFN (Balazadeh et al., 2025) and (8) DoPFN (Robertson et al., 2025). Further implementation details are in Appendix D.3.
Results on standard CATE estimation. We benchmark our model on ten synthetic datasets generated under diverse mechanisms, with implementation details in Appendix D. In addition, we evaluate on a semi-synthetic version of the Jobs dataset (Smith and Todd, 2005), derived from the widely used LaLonde study (LaLonde, 1986). Here, we generate outcomes to create a semi-synthetic dataset and allow for evaluation against ground-truth.
@@ 6.1 Evaluation for standard CATE setting
MRIV (Frauen and Feuerriegel, 2022) 0.688 $\pm$ 0.21 0.641 $\pm$ 0.24
@@
Baselines for IV setting. We benchmark against a broad set of state-of-the-art IV methods for treatment effect estimation: (1) KIV (Singh et al., 2019): a nonlinear extension of two-stage least squares using kernel ridge regression with feature maps; (2) DFIV (Xu et al., 2021): extends KIV by parameterizing feature maps with neural networks trained iteratively; (3) DeepIV (Hartford et al., 2017): a two-stage neural approach, first estimating the treatment distribution and then solving a counterfactual prediction task; (4) DeepGMM (Bennett et al., 2019): formulates IV estimation as a minimax game based on the generalized method of moments, solved via adversarial training; (5) DMLIV (Syrgkanis et al., 2019): a double machine learning framework that estimates nuisance functions and learns the CATE by orthogonalized regression; (6) DRIV (Syrgkanis et al., 2019): a meta-learner combining DMLIV with doubly robust pseudo-outcomes for improved stability; and (7) MRIV (Frauen and Feuerriegel, 2022): a multiply robust framework for binary IVs that directly estimates CATE via pseudo–outcome regression. For foundation model baselines, as CausalPFN (Balazadeh et al., 2025) is only for back-door adjustment, we include DoPFN (Robertson et al., 2025).
@@ 6.3 Front-door adjustment
We additionally evaluate our model under the front-door adjustment setting in Table 6.2. Due to space constraints, details are provided in Appendix H.1. The experiments show the flexibility of our method to perform causal inference in the front-door adjustment setting.
6.4 Discussion
Limitations and future work. The current evaluation is limited to synthetic and semi-synthetic data due to the fundamental problem of missing potential outcomes on real-world data. For future work, it will be interesting to investigate the performance of CausalFM in applied A/B experimental setups to assess its empirical performance and robustness under real-world conditions. Additionally, an important research direction will be to incorporate interpretability or fairness constraints into CausalFM, which is crucial for reliable deployment in practice.
Ethics statement.
Human subjects and IRB. This work does not involve experiments with human subjects. Our training data are synthetically generated from prespecified SCM-based priors. For empirical evaluation, we additionally use publicly available benchmark data (e.g., Jobs) where outcomes are generated in a semi-synthetic manner following common practice; no identifiable personal information is introduced by us. Accordingly, no IRB review was required for this study.
Data, privacy, and security. We do not collect, store, or release sensitive personal data. Public datasets are used under their respective licenses, and our semi-synthetic outcome generation avoids re-identification risks. We will document preprocessing and generation steps to support reproducibility.
Bias and potential harms. Causal estimators can be misused if applied outside the assumed identification regime (e.g., back-door, front-door, IV) or under severe violations (e.g., weak instruments, lack of overlap). To mitigate harm: (i) we make assumptions explicit and provide uncertainty quantification; (ii) we advocate domain-expert validation and sensitivity checks before deployment; (iii) we discourage high-stakes automated decision-making without human oversight.
Use of large language models (LLMs). We used LLM-based tools to assist with writing (clarity, grammar) and for literature research. All claims were authored and verified by the authors; citations were cross-checked against primary sources. No sensitive data were provided to LLM tools.
Reproducibility statement. We ensure reproducibility of our results by providing the full implementation and training scripts. Our CausalFM Toolkit is available at https://github.com/yccm/CausalFM-toolkit; Toolkit docs are at https://causalfm-toolkit.readthedocs.io; Project code is available at https://github.com/yccm/CausalFM. The repository contains the necessary code to reproduce our experiments, along with instructions for dataset preparation, model training and evaluation procedures. This setup allows independent researchers to replicate the reported results and extend our work with minimal effort.
Acknowledgments. This paper is supported by the DAAD program “Konrad Zuse Schools of Excellence in Artificial Intelligence”, sponsored by the Federal Ministry of Research, Technology, and Space. Additionally, this work has been supported by the German Federal Ministry of Education and Research (Grant: 01IS24082).